Find the equilibrium of L and K subject to a given output:
Minimize cost: C = 40K+ 20L Subject to Q = 60 L K= 3000
Sir please solve this question
1
Expert's answer
2018-02-21T09:43:07-0500
1) So, L*K = 3000, L = 3000/K;
2) C = 40K + 20L, for minimizing cost we should get that MCK = 40, MCL = 20;
3) L/K = MCL / MCK = 20/40 = 0.5, L = 0.5*K
0.5*K = 3000/K, K = 77.5, L = 38.7
C = 40 * 77.5 + 20 * 38.7 = 3874